OpenStudy (anonymous):

A company builds computers. It costs $7,600 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? C(x) = 460x − 3000 C(x) = 3000x − 460 C(x) = 460x + 3000 C(x) = 3000x + 460

OpenStudy (pfenn1):

Let x = number of computers built, A = a constant, B = unit cost to build a computer. \[7600=A+10B\]\[12200=A+20B\]

OpenStudy (pfenn1):

You have two equations and two unknowns. Do you know how to solve for A and B?